OpenStudy (anonymous):

let finction g be defined by g(x)=x-2. If 4g(y)=18, what is the value of g(3y)? some one save meee, how do you solve this?

OpenStudy (anonymous):

4g(y)=18 g(y) = 18/4 = 4.5

OpenStudy (anonymous):

g(x) = x-2 g(y) = y - 2 g(y) = 4.5 4.5 = y-2 y = 2 + 4.5 = 6.5

OpenStudy (anonymous):

g(x) = x-2 g(3y) = 3y-2 = 6.5*3 -2 = ?

OpenStudy (anonymous):

thank you so much! mind helping me on my other questions? :))